# Predicting the Geolocation of Tweets Using transformer models on Customized Data

### v3

@@ Appendix A Per-user geolocation estimation
Appendix B Loss functions
Figure 8: Multiple Outcomes Prediction (MOP) model loss functions computational graph including visualization of Weighted Linear Combination (WLC) and SoftMax (SM) components
Figure 9: Total loss functions computational graph for the MOP-type Key Feature (KF) with M outcomes, and K Minor Features (MF) of the SOP-type
Appendix C Performance metrics
C.0.1 Geospatial metrics
  • SAE - simple accuracy error metric is utilized to measure the geographic distance between two points on the Earth’s surface. This is achieved through the application of the Haversine formula for calculating the great-circle distance in kilometers. The formula is as follows:
@@ C.0.1 Geospatial metrics
  • Acc@161 - the percentage of predicted locations that are within a 161km (100 miles) radius of the actual location.
@@
C.0.2 Probabilistic metrics
  • CAE - comprehensive accuracy error metric measures the expected distance between the true origin of the tweet and a random point generated from the GMM predicted by the model. The Haversine formula is used to calculate distances from the original point to each Gaussian sub-population of 100 samples. Bakerman in his work [2] suggests the Monte Carlo method to compute CAE as the mean of the sub-population distance errors:

### v4

@@ Appendix A Per-user geolocation estimation
Appendix B Performance metrics
B.1 Geospatial metrics
  • SAE—simple accuracy error metric is utilized to measure the geographic distance between two points on the Earth’s surface. This is achieved through the application of the Haversine formula for calculating the great-circle distance in kilometers. The formula is as follows:
@@ B.1 Geospatial metrics
  • Acc@161—the percentage of predicted locations that are within a 161km (100 miles) radius of the actual location.
@@
B.2 Probabilistic metrics
  • CAE—comprehensive accuracy error metric measures the expected distance between the true origin of the tweet and a random point generated from the GMM predicted by the model. The Haversine formula is used to calculate distances from the original point to each Gaussian sub-population of 100 samples. Bakerman in his work [2] suggests the Monte Carlo method to compute CAE as the mean of the sub-population distance errors:
@@ B.2 Probabilistic metrics
Appendix C Loss functions
Figure 8: Multiple Outcomes Prediction (MOP) loss functions computational graph including visualization of Weighted Linear Combination (WLC) and SoftMax (SM) components.
$\widehat{Y}_{m}$—point, $\widehat{c}_{m}$—covariance parameter, and $\widehat{w}_{m}$—weight predicted for a single outcome $m$
$D_{m}^{2}$—squared error distance, $\sigma_{m}$—adjusted covariance (confidence), and $\widehat{W}_{m}$—normalized weight (significance) for a single outcome $m$
Figure 9: Total loss functions computational graph for the MOP-type Key Feature (KF) with M outcomes, and K Minor Features (MF) of the SOP-type.